OpenStudy (anonymous):

Two angles are supplementary if the sum of their measures is 180 degrees. The measure of the first angle is 54 degrees more than three times the second angle. Find the measure of the first angle.

OpenStudy (lgbasallote):

let x = first angle y = 2nd angle x+ y = 180 x = 54 +3y 54 + 3y + y = 180 54 + 4y = 180 4y = 126 y= 31.5 x = 148.5 the first angle is 148.5 degrees